two parallels wires, wire P and Q which carries a current of 30A and 50A, respectively in the same direction.The wires are placed 1.25 mm apart in a vacuum.



a) Calculate the magnetic force per unit length on wire P.



b) Determine the position x from wire P in between the wires where the magnetic field is zero.

1
Expert's answer
2022-02-23T08:30:49-0500

a)

"F= \n\\frac{\n\u03bc _\n0\n\u200b\n I _\n1\n\u200b\n I _\n2}{2\\pi d}\n\u200b\n \n\u200b\n =240 mN,"

b)

"x=\frac{dI_1}{I_2}=0.75~mm."